. There are 250 dogs at a dog show that weigh an average of 12 pounds, with

a standard deviation of 8 pounds. If 4 dogs are chosen at random, what is

the probability that the average weight is greater than 8 pounds?


1
Expert's answer
2021-06-02T10:13:26-0400

"\mu=12 \\

\sigma=8 \\

n=4 \\

P(\bar{X}>8) = 1 -P(\bar{X}<8) \\

= 1 -P(Z< \frac{8-12}{8 / \sqrt{4}}) \\

= 1 -P(Z < -1) \\

= 1 -0.1586 \\

= 0.8414"